2019 LBCs annual maintenance

Summary

Date: August 5-9, 2019

Location: Electronic Lab Room 116

Performed by: Fernando Pedichini & Alessandro Cavallaro.

  • A general clean up of the cameras from dust and dirt was carried out. The shutter was gently cleaned with some light ethanol.
  • Checked proper operation of both shutters.
  • A visual inspection of the cryostat windows was performed (by keeping the shutter open) for both LBCs in order to check for evidence of damage and for mylar residues. Nothing of concern was noticed, only the already known coating delamination on LBC Blue L6 (see IT#4845): Inspection reveals LBC Red clock cable not properly connected:

The cable was fixed before to proceed with the tests.

As for last year, an old version of IDL was installed in the Lab PC, so the date of the computer had to be changed to 2015 to make the IDL license work. Therefore the FITS file names are off by 3 years (month and day are OK).

Fernando wrote an IDL script to capture and analyze a sequence of images, recording the bias and RON histograms and the bias and RON evolution in time. Screenshots of the graphs were recorded.

We proceeded with warm tests followed by cold test for both cameras (using LN2 to cool down the cameras). The vacuum in the cryostats was a little bit soft so we pumped them before and during cool down.

Vacuum

Before [mBar]

After [mBar]

LBC Red

n/a

1.30 e-05

LBC Blue

5.05 e-04

2.02 e-05

An all-night-long test revealed 4 images with bias problem on controller#2 (see IT#7750).

The morning after LBC Red tech controller (controller #2) suddenly stopped responding to any command. We did not find any smoking gun so we decided to replace it with Spare#3 and send controller#2 to Skytech.

Fernando and I tried to take care of the open ITs for the LBCs. Here is the initial and final situation:

LBCs IssueTraks

Before

After

Critical

0

0

High

6

1

Medium

40

34

Low

8

5

Monitor

7

8

All the remaining ITs are not electrical nor mechanical, except for a couple of them used to track documentation and except IT#7737 regarding the LBC Red rotator limit switch. The remaining ITs regard software and image quality. Next action can be for Alessandro to meet John Hill and Olga Kuhn to discuss them.

RED SCIENCE LOG

  • Controller#1
  • Vacuum pressure: 1.30 e-05
  • Temp controller: 40C
  • Temp CCD: 179K
  • Reference images (full FITS files):
    • ch1: lbcRsci.20160107.162025_1.fits
    • ch2: lbcRsci.20160107.162025_2.fits
    • ch3: lbcRsci.20160107.162025_3.fits
    • ch4: lbcRsci.20160107.162025_4.fits

Channel

1

2

3

4

Housekeeping OFF

Bias level (ADUs)

299

303

297

299

RON (ADUs)

5.0

4.8

5.0

4.7

Offset (V)

2.5870

2.6103

2.6148

2.5899

Housekeeping ON

Bias level (ADUs)

296

302

288

289

RON (ADUs)

6.4

4.9

5.1

5.2

Power Supplies

Supply

Voltage measured (V)

Telemetry (V)

+8 VD

9.69

-

+5 VD

5.07

4.55

+11 VA

11.01

10.98

-11 VA

-10.94

-11.00

+25 VA

25.06

24.96

-25 VA

-24.82

-24.85

RED TECH LOG

  • Vacuum pressure: 1.30 e-05
  • Temp controller: 44C
  • Temp CCD: 179K
  • Reference images (full FITS files):
    • ch1: lbcRtec.20160107.145536_1.fits
    • ch2: lbcRtec.20160107.145536_2.fits
Housekeeping OFF

Channel

1

2

Bias level (ADUs)

298

298

RON (ADUs)

5.7

5.4

Offset (V)

2.5478

2.5664

BLUE SCIENCE LOG

  • Controller #6
  • Vacuum pressure: 2.02 e-05
  • Temp controller: 60C
  • Temp CCD: 205K
  • Reference images (full FITS files):
    • ch1: lbcBsci.20160107.101046_1.fits
    • ch2: lbcBsci.20160107.101046_2.fits
    • ch3: lbcBsci.20160107.101046_3.fits
    • ch4: lbcBsci.20160107.101046_4.fits

Channel

1

2

3

4

Housekeeping OFF

Bias level (ADUs)

305

305

304

305

RON (ADUs)

4.5

4.7

4.6

4.8

Offset (V)

2.5602

2.5626

2.5710

2.5651

Housekeeping ON

Bias level (ADUs)

291

291

293

292

RON (ADUs)

4.8

5.3

5.3

5.3

Power Supplies

Supply

Voltage measured (V)

Telemetry (V)

+8 VD

8.92

-

+5 VD

5.31

4.92

+11 VA

10.94

10.77

-11 VA

-11.00

-11.02

+25 VA

25.02

24.82

-25 VA

-24.96

-24.90

BLUE TECH

  • Controller#: 4
  • Vacuum pressure: 2.02 e-05
  • Temp controller: 64C
  • Reference images (full FITS files):
    • ch1: lbcBtec.20160107.082649_1.fits
    • ch2: lbcBtec.20160107.082649_2.fits
Housekeeping OFF

Channel

1

2

Bias level (ADUs)

299

299

RON (ADUs)

5.9

7.1

Offset (V)

2.5350

2.5730

The tangential fan seems to be a little slow. There’s no feedback on the speed, maybe a simple system to measure it can be done / purchased in the future.

Note: the images have been taken with the lights ON, we noticed a little difference between bias and dark.

All the relevant images taken during the 2019 maintenance are on the LBC’s computer, in the folder on the desktop named “2019 LBC’s annual maintenance”

-- AlessandroCavallaro - 10 Sep 2019
Topic revision: r1 - 10 Sep 2019, AlessandroCavallaro
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback